# From Sim-to-Real: Toward General Event-based Low-light Frame Interpolation with Per-scene Optimization

### v1

@@ Abstract
Video Frame Interpolation (VFI) is important for video enhancement, frame rate up-conversion, and slow-motion generation. The introduction of event cameras, which capture per-pixel brightness changes asynchronously, has significantly enhanced VFI capabilities, particularly for high-speed, nonlinear motions. However, these event-based methods encounter challenges in low-light conditions, notably trailing artifacts and signal latency, which hinder their direct applicability and generalization. Addressing these issues, we propose a novel per-scene optimization strategy tailored for low-light conditions. This approach utilizes the internal statistics of a sequence to handle degraded event data under low-light conditions, improving the generalizability to different lighting and camera settings. To evaluate its robustness in low-light condition, we further introduce EVFI-LL, a unique RGB+Event dataset captured under low-light conditions. Our results demonstrate state-of-the-art performance in low-light environments. Both the dataset and the source code will be made publicly available upon publication. Project page: https://naturezhanghn.github.io/sim2real.
Figure 1: The interpolated result from the real-captured RGB-Event sequence under low light conditions. Our proposed per-scene optimization method can successfully correct the impact of event latency, accurately interpolate the correct positions, and produce visually pleasing interpolation results.
$\ast$ Authors contributed equally $\dagger$ Corresponding author
1 Introduction
Video frame interpolation (VFI) has extensive applications in video enhancement, frame rate up-conversion, and slow-motion generation. The advent of event cameras [24, 2] has demonstrated a good potential for VFI, because it can asynchronously capture of per-pixel brightness changes with much higher temporal resolution and higher dynamic range. Recent studies [30, 29, 28, 11] have explored the event-based video frame interpolation (Event-VFI), demonstrating significant performance enhancements compared to the traditional RGB-based VFI [1, 26, 7].
However, the performance of Event-VFI degrades in low light captures due to the noticeable trailing artifacts associated with event signals [17, 15, 20]. As shown in Fig. 1 (d) and (e), one can observe that in the results from the current Event-VFI methods, such as TimeLens [30] and CBMNet [11], there is noticeable ghosting of letters due to event latency, and the interpolated frames exhibit noticeable trailing artifacts. Given that low light conditions are common, investigating Event-VFI under low light conditions represents a meaningful and practical yet challenging task.
The low-light condition presents another challenge to Event-VFI: the latency, noise, and other degradations in the event signals are not easy to model. One line of solution [14, 34, 13] is to simulate more realistic events by accounting for various degradation factors, including leak noise events, refractory periods, and shot noise, using event simulators, such as v2e [6] or ESIM [21]. However, even with an improved simulation pipeline, a significant discrepancy between simulated data and real-world data still exists [17]. Another solution [17] focuses on collecting real-world event datasets to bridge the gap between simulated and actual event signals. However, dataset acquisition is costly, and the datasets tailored to one particular camera with a fixed setting cannot generalize to other cameras or even to other settings. Thus, this raises an interesting and challenging question: how can the Event-VFI method be cost-effectively generalized to real-world low-light conditions with varying camera settings?
To address the above question, we propose a novel per-scene optimization strategy for general Event-VFI in low light conditions. Because the trailing artifacts and other degradations in event signals are hard to model at the training stage, we instead perform per-scene optimization to adapt an interpolation model to a specific distribution of that event stream. This is based on the insight that these degradations are mostly influenced by illuminance, camera hardware, and capture settings, which are relatively constant in a single video. As shown in Fig. 2, unlike previous Event-VFI methods [30, 29, 11], our method first utilizes the entire input sequence to fine-tune a pre-trained Event-VFI model. Specifically, we use the captured RGB image $I_{t}$ at time $t$ as the ground truth (with denoising applied) to supervise the interpolation results from two other captured RGB frames ($I_{t-1}$ and $I_{t+1}$) at neighboring timestamps and event signals ($E_{t-1,t}$ and $E_{t,t+1}$) during this interval. Then, at the actual inference stage, the fine-tuned network is applied to the same sequence to generate a novel frame $\hat{I}_{t+\Delta t}$ to actually increase the frame rate.
Figure 2: Illustration of the proposed per-scene optimization process.
Additionally, we also improve the robustness of pre-trained Event-VFI in low-light conditions. We revisit and refine the low-light event modeling process [20, 17, 5, 22, 4], carefully designing a simulation pipeline tailored for low-light Event-VFI. Utilizing our pre-trained model, our per-scene optimized approach achieves state-of-the-art performance in processing real-world low-light sequences. Additionally, the running time for per-scene optimization is comparable to that of the inference process, thereby ensuring that these improvements are achieved at minimal cost.
To evaluate the proposed solution, we also developed EVFI-LL, a novel RGB-Event interpolation benchmark designed for low-light. Unlike previous datasets captured at normal lighting conditions, such as BS-ERGB [30] and ERF-X170FPS [11], EVFI-LL consists of sequences captured under low-light conditions under 35 Lux. Notably, EVFI-LL is the first dataset to capture events using different ON/OFF thresholds (three settings: -20, 0 [default], and 20) on the Prophesee EVK4-HD event sensor. Comprising over 20 testing sequences, this dataset includes accurately synchronized RGB videos and event streams, establishing a new benchmark for evaluating low-light Event-VFI. Additionally, it allows for an assessment of the generalizability of Event-VFI methods across various ON/OFF thresholds. Both the code and dataset will be made publicly available upon publication.
Figure 3: Overview of the proposed low light Event-VFI method. This framework encompasses the pre-training phase, per-scene optimization, and inference process.
2 Related Work
2.1 RGB-based Video Frame Interpolation
Video frame interpolation is a fundamental and challenging task that aims to generate inter-frames between consecutive frames in a video. Typical solutions for RGB-based video frame interpolation (RGB-VFI) can be classified into flow-free and flow-based methods. For the flow-free methods, the interpolated frame is generated directly from input frames without explicit flow estimation  [19, 18, 25]. [19] utilized adaptive separable convolution for interpolation, achieved trained end-to-end using widely available video data without any human annotation. PhaseNet  [18] estimated the phase decomposition of the intermediate frame. [25] proposed a Transformer-based interpolation strategy that reduces memory consumption and improves efficiency through local attention and a space-time separation strategy.
For the Flow-based methods [1, 26, 7], the flow is explicitly estimated in the interpolation. [1] developed a depth-aware flow projection layer to synthesize intermediate flows that preferably sample closer objects than farther ones. [26] proposed bidirectional optical flow learning for interpolating large motions and complex textures. [7] introduced RIFE, which can estimate the intermediate flows end-to-end with much faster speed. However, due to a lack of accurate motion information between two RGB frames, the performance of RGB-VFI degrades when handling complex and large motion.
@@ 2.1 RGB-based Video Frame Interpolation
2.2 Event-based Video Frame Interpolation
@@
Event cameras record intensity changes of objects, offering high frame rate and high dynamic range characteristics [15, 20], which is benefit for VFI task. Thus, using event cameras for VFI (Event-VFI) has recently provided a new direction for addressing non-linear, large motion interpolation [30, 29, 28, 11]. Time Lens [30] and Time Lens++ [29] demonstrated promising performance in non-linear motion scenarios. [28] proposed REFID, which jointly performs image deblurring and interpolation. [11] introduced CBMNet for interpolation with cross-modal asymmetric bidirectional motion field estimation.
@@
Even though Event-VFI methods exhibit better performance in handling non-linear motions, current research has not addressed Event-VFI task in low-light conditions. In low light, event cameras produce noticeable trailing artifacts [17, 15, 20], which negatively impact event-based interpolation. Thus in this paper, we study this promising yet under-explored low light Event-VFI problem.
2.3 Low-light Event Correction
@@
In low light, event cameras produce noticeable trailing artifacts and noise [17, 15, 20] which degrade the follow-up tasks. To correct these degradation in low-light events, some methods [14, 34] utilize more realistic event simulation. Specially, [14, 34] utilized the commonly used v2e [6] or ESIM [21] simulator, and considered various degradation factors such as threshold noise, hot pixels, leak noise events, refractory period, and shot noise. However, event signal generated by such simulation method still has great gap with real-captured data and the model trained with simulated data can not directly generalize to real-world events and correct the trailing artifacts of low light events [17].
@@
To correct the real-world latency and noise of event signal, [17] collected a real-world low-light event dataset consisting of low-light events and high-quality images. A learnable event timestamps calibration model is also proposed to learn the event trailing. However, such method needs costly data collection. More importantly, such methods [17] mainly focusing on correct event signal for specific event camera, specifically for Prophesee EVK4, with fixed camera setting, e.g., ON/OFF thresholds. However, the hardware differences and parametric setting both effect the latency degree, limiting their generalization. To solve these limitations, in this paper, a per-scene optimization strategy is proposed for low light Event-VFI task to let our model can generalize to event camera with different camera setting without costly data collection.
3 Problem definition
3.1 Low-light Event Representation
@@
Event camera capture scene illumination change as an asynchronous stream of events. Each event at position $\mathbf{u}=(x,y)$ is generated when the intensity change is larger than a contrast threshold $c$[30, 29, 11]:
$E_{t}=\begin{cases}1,&\text{if }\log(I_{t})-\log(I_{t-\Delta t})\geq c,\\ -1,&\text{if }\log(I_{t})-\log(I_{t-\Delta t})\leq-c,\\ 0,&\text{otherwise},\end{cases}$ (1)
@@
However, event streams in low light often contain trailing artifacts [17, 15, 20] (Fig. 4(d)), which is not covered by the simplified modeling in Eqn. 1. To better model it, we revisit event generation process. As shown in Fig. 4 (a), the photodiode (PD) converts an incoming incident light to photocurrent, whose magnitude is effected by the cutoff frequency $f_{c}$ of the camera system [20, 17] as
$f_{c}=\frac{1}{2\pi C}\frac{I_{ph}}{U_{t}},$ (2)
@@
$I^{\prime}_{t}=\alpha I_{t}+(1-\alpha)I_{t-\Delta t},$ (3)
@@
$E_{t}=\begin{cases}1,&\text{if }\log(\alpha I_{t}+(1-\alpha)I_{t-\Delta t})-\log(I_{t-\Delta t})\geq c,\\ -1,&\text{if }\log(\alpha I_{t}+(1-\alpha)I_{t-\Delta t})-\log(I_{t-\Delta t})\leq-c,\\ 0,&\text{otherwise}.\end{cases}$ (4)
Figure 4: Simplified circuit diagram for each pixel of the event camera. And the visualization of event trailing artifacts, captured by our EVS-RGB beam splitting imaging system.
Equations above explain the cause of trailing. When image intensity decreases, the cutoff frequency also decreases (Eqn. 2), leading to a smaller value of $\alpha$. Consequently, the measurement of current intensity is more significantly influenced by prior values, an effect modeled as an RC low-pass filter [6, 17]. Such a configuration implies that the same voltage changes can induce more extended temporal delays in the generation of event signals in the low light condition. In Fig. 4(d), we display a real-world captured low-light event, where a temporal delay in the event data can be observed, leading to spatial trailing artifacts. Using such degraded event data for frame interpolation results in deformations in the interpolated frames due to the trailing effects (see Fig. 1).
3.2 Low Light Event-VFI
For common Event-VFI, the interpolation is modeled using Eqn. 1 and formulated as [9, 27, 16]:
@@ 3.2 Low Light Event-VFI
In recent methods [30, 29, 11], to recover an intermediate RGB frame $I_{t+\Delta t}$ between $I_{t}$ and $I_{t+1}$, information from both the forward (from $I_{t}$ to $I_{t+\Delta t}$) and the backward (from $I_{t+1}$ to $I_{t+\Delta t}$) estimation are utilized. Thus process of low light Event-VFI in Eqn. 6 is solved as:
$\hat{I}_{t+\Delta t}=\Phi(I_{t},I_{t+1},E_{t,t+\Delta t},E_{t+\Delta t,t+1},c,\alpha),$ (7)
where $\Phi$ is the interpolation.

### v2

@@ Abstract.
Video Frame Interpolation (VFI) is important for video enhancement, frame rate up-conversion, and slow-motion generation. The introduction of event cameras, which capture per-pixel brightness changes asynchronously, has significantly enhanced VFI capabilities, particularly for high-speed, nonlinear motions. However, these event-based methods encounter challenges in low-light conditions, notably trailing artifacts and signal latency, which hinder their direct applicability and generalization. Addressing these issues, we propose a novel per-scene optimization strategy tailored for low-light conditions. This approach utilizes the internal statistics of a sequence to handle degraded event data under low-light conditions, improving the generalizability to different lighting and camera settings. To evaluate its robustness in low-light condition, we further introduce EVFI-LL, a unique RGB+Event dataset captured under low-light conditions. Our results demonstrate state-of-the-art performance in low-light environments. Project page: https://openimaginglab.github.io/Sim2Real/.
@@
Event Camera, Video Frame Interpolation, Low Light, Per-scene Optimization
Figure 1. The interpolated result from the real-captured RGB-Event sequence under low light conditions. Our proposed per-scene optimization method can successfully correct the impact of event latency, accurately interpolate the correct positions, and produce visually pleasing interpolation results.
1. Introduction
High-speed imaging is important to computational photography and computer graphics, facilitating the capture of rapidly moving objects and scenes (Yang et al., 2022; Zhang et al., 2021; Jiang et al., 2018; Choi and Kweon, 2020). This is also critical in other complex applications, such as turbulence visualization (Bai et al., 2021; Wang et al., 2024), fast fluorescence lifetime imaging (Lee et al., 2019), and high-frequency vibration analysis (Liu et al., 2005). Although high-speed cameras are a widely used solution, their high cost and need for intense illumination often restrict their practicality. An effective alternative is video frame interpolation (VFI), which generates high-frame-rate videos from low-frame-rate videos captured with conventional cameras. VFI has been widely applied in video enhancement (Choi and Kweon, 2020) and slow-motion generation (Jiang et al., 2018). The advent of event cameras (Lichtsteiner et al., 2008; Brandli et al., 2014; Guo et al., 2023; Kodama et al., 2023) has demonstrated good potential for VFI, because it can asynchronously capture per-pixel brightness changes with much higher temporal resolution and higher dynamic range. Recent studies (Tulyakov et al., 2021, 2022; Sun et al., 2023; Kim et al., 2023) have explored the event-based video frame interpolation (Event-VFI), demonstrating significant performance enhancements compared to the traditional RGB-based VFI (Bao et al., 2019; Sim et al., 2021; Huang et al., 2022).
However, the performance of Event-VFI degrades in low light captures due to the noticeable trailing artifacts associated with event signals (Liu et al., 2024; Graça et al., 2023; Graca and Delbruck, 2021; Hu et al., 2021). As shown in Fig. 1 (d) and (e), one can observe that in the results from the current Event-VFI methods, such as TimeLens (Tulyakov et al., 2021) and CBMNet (Kim et al., 2023), there is ghosting of letters due to event latency, and the interpolated frames exhibit noticeable trailing artifacts. Given that low-light conditions are common, investigating Event-VFI under low-light conditions represents a meaningful, practical, yet challenging task.
The low-light condition presents another challenge to Event-VFI: the latency, noise, and other degradations in the event signals are difficult to model. One line of solution (Liang et al., 2023; Zhou et al., 2023; Liang et al., 2024) is to simulate more realistic events by accounting for various degradation factors, including leak noise events, refractory periods, and shot noise, using event simulators, such as v2e (Hu et al., 2021) or ESIM (Rebecq et al., 2018). However, even with an improved simulation pipeline, a significant discrepancy between simulated data and real-world data still exists (Liu et al., 2024). Another solution (Liu et al., 2024) focuses on collecting real-world event datasets to bridge the gap between simulated and actual event signals. However, dataset acquisition is costly, and the datasets tailored to one particular camera with a fixed setting cannot generalize to other cameras or even to other settings. Thus, this raises an interesting and challenging question: how can the Event-VFI method be cost-effectively generalized to real-world low-light conditions with varying camera settings?
To address the above question, we propose a novel per-scene optimization strategy for general Event-VFI in low light conditions. Because the trailing artifacts and other degradations in event signals are hard to model at the training stage, we instead perform per-scene optimization to adapt an interpolation model to a specific distribution of that event stream. This is based on the insight that these degradations are mostly influenced by illuminance, camera hardware, and capture settings, which are relatively constant in a single video. As shown in Fig. 2, unlike previous Event-VFI methods (Tulyakov et al., 2021, 2022; Kim et al., 2023), our method first utilizes the entire input sequence to fine-tune a pre-trained Event-VFI model. Specifically, we use the captured RGB image $I_{t}$ at time $t$ as the ground truth (with denoising applied) to supervise the interpolation results from two other captured RGB frames ($I_{t-1}$ and $I_{t+1}$) at neighboring timestamps and event signals ($E_{t-1,t}$ and $E_{t,t+1}$) during this interval. Then, at the actual inference stage, the fine-tuned network is applied to the same sequence to generate a novel frame $\hat{I}_{t+\Delta t}$ to actually increase the frame rate.
Figure 2. Illustration of the proposed per-scene optimization process.
Additionally, we also improve the robustness of pre-trained Event-VFI in low-light conditions. We revisit and refine the low-light event modeling process (Posch et al., 2010; Liu et al., 2024; Graça et al., 2023; Rios-Navarro et al., 2023; Graca and Delbruck, 2021), carefully designing a simulation pipeline tailored for low-light Event-VFI. Utilizing our pre-trained model, our per-scene optimized approach achieves state-of-the-art performance in processing real-world low-light sequences. Additionally, the running time for per-scene optimization is comparable to that of the inference process, thereby ensuring that these improvements are achieved at minimal cost.
To evaluate the proposed solution, we also developed EVFI-LL, a novel RGB-Event interpolation benchmark designed for low-light. Unlike previous datasets captured at normal lighting conditions, such as BS-ERGB (Tulyakov et al., 2021) and ERF-X170FPS (Kim et al., 2023), EVFI-LL consists of sequences captured under low-light conditions under 35 Lux. Notably, EVFI-LL is the first dataset to capture events using different ON/OFF thresholds (three settings: -20, 0 [default], and 20) on the Prophesee EVK4-HD event sensor. Comprising over 20 testing sequences, this dataset includes accurately synchronized RGB videos and event streams, establishing a new benchmark for evaluating low-light Event-VFI. Additionally, it allows for an assessment of the generalizability of Event-VFI methods across various ON/OFF thresholds.
Figure 3. Overview of the proposed low light Event-VFI method. This framework contains the pre-training phase, per-scene optimization, and inference process.
2. Related Work
2.1. RGB-based Video Frame Interpolation
Video frame interpolation is a fundamental and challenging task that aims to generate inter-frames between consecutive frames in a video. Typical solutions for RGB-based video frame interpolation (RGB-VFI) can be classified into flow-free and flow-based methods. For the flow-free methods, the interpolated frame is generated directly from input frames without explicit flow estimation, with adaptive separable convolution (Niklaus et al., 2017), phase decomposition (Meyer et al., 2018) and transformer-based structure (Shi et al., 2022). For the Flow-based methods (Bao et al., 2019; Sim et al., 2021; Huang et al., 2022; Wu et al., 2022), the optical flow is explicitly estimated in the interpolation. However, due to a lack of accurate motion information between two RGB frames, the performance of RGB-VFI degrades when handling complex and large motion.
2.2. Event-based Video Frame Interpolation
@@
Event cameras record intensity changes of objects, offering high frame rate and high dynamic range characteristics (Posch et al., 2010; Chen et al., 2024; Gehrig and Scaramuzza, 2024), which is benefit for VFI task. Thus, using event cameras for VFI (Event-VFI) has recently provided a new direction for addressing non-linear, large motion interpolation (Tulyakov et al., 2021, 2022; Sun et al., 2023; Kim et al., 2023). Time Lens (Tulyakov et al., 2021) and Time Lens++ (Tulyakov et al., 2022) demonstrated promising performance in non-linear motion scenarios. Sun et al. (2023) proposed REFID, which jointly performs image deblurring and interpolation. Kim et al. (2023) introduced CBMNet for interpolation with cross-modal asymmetric bidirectional motion field estimation.
@@
Even though Event-VFI methods exhibit better performance in handling non-linear motions, current research has not addressed Event-VFI task in low-light conditions. In low light, event cameras produce noticeable trailing artifacts (Liu et al., 2024; Posch et al., 2010), which negatively impact event-based interpolation. Thus in this paper, we study this promising yet under-explored low light Event-VFI problem.
2.3. Low-light Event Correction
@@
In low light, event cameras produce noticeable trailing artifacts and noise (Liu et al., 2024; Posch et al., 2010; Hu et al., 2021) which degrade the follow-up tasks. To correct these degradation in low-light events, some methods (Liang et al., 2023; Zhou et al., 2023) utilize more realistic event simulation. Specially, Liang et al. (2023); Zhou et al. (2023) utilized the commonly used v2e (Hu et al., 2021) or ESIM (Rebecq et al., 2018) simulator, and considered various degradation factors such as threshold noise, hot pixels, leak noise events, refractory period, and shot noise. However, event signal generated by such simulation method still has great gap with real-captured data and the model trained with simulated data can not directly generalize to real-world events and correct the trailing artifacts of low light events (Liu et al., 2024).
@@
To correct the real-world latency and noise of event signal, Liu et al. (2024) collected a real-world low-light event dataset consisting of low-light events and high-quality images. A learnable event timestamps calibration model is also proposed to learn the event trailing. However, such method needs costly data collection. More importantly, such methods (Liu et al., 2024) mainly focusing on correct event signal for specific event camera, specifically for Prophesee EVK4, with fixed camera setting, e.g., ON/OFF thresholds. However, the hardware differences and parametric setting both effect the latency degree, limiting their generalization. To solve these limitations, in this paper, a per-scene optimization strategy is proposed for low light Event-VFI task to let our model can generalize to event camera with different camera setting without costly data collection.
3. Problem definition
3.1. Low-light Event Representation
@@
Event camera capture scene illumination change as an asynchronous stream of events. Each event at position $\mathbf{u}=(x,y)$ is generated when the relative intensity change is larger than a contrast threshold $c$ (Tulyakov et al., 2021, 2022; Kim et al., 2023):
(1) $E_{t}=\begin{cases}1,&\text{if }\log(I_{t})-\log(I_{t-\Delta t})\geq c,\\ -1,&\text{if }\log(I_{t})-\log(I_{t-\Delta t})\leq-c,\\ 0,&\text{otherwise},\end{cases}$
@@
However, event streams in low light often contain trailing artifacts (Hu et al., 2021; Liu et al., 2024; Posch et al., 2010) (Fig. 4(d)), which is not covered by the simplified modeling in Eqn. 1. To better model it, we revisit event generation process. As shown in Fig. 4 (a), the photodiode (PD) converts an incoming incident light to photocurrent, whose magnitude is effected by the cutoff frequency $f_{c}$ of the camera system (Posch et al., 2010; Liu et al., 2024) as
(2) $f_{c}=\frac{1}{2\pi C}\frac{I_{ph}}{U_{t}},$
@@
(3) $I^{\prime}_{t}=\alpha I_{t}+(1-\alpha)I_{t-\Delta t},$
@@
(4) $E_{t}=\begin{cases}1,&\text{if }\log(\alpha I_{t}+(1-\alpha)I_{t-\Delta t})-\log(I_{t-\Delta t})\geq c,\\ -1,&\text{if }\log(\alpha I_{t}+(1-\alpha)I_{t-\Delta t})-\log(I_{t-\Delta t})\leq-c,\\ 0,&\text{otherwise}.\end{cases}$
Equations above explain the cause of trailing. When image intensity decreases, the cutoff frequency also decreases (Eqn. 2), leading to a smaller value of $\alpha$. Consequently, the measurement of current intensity is more significantly influenced by prior values, an effect modeled as an RC low-pass filter (Hu et al., 2021; Liu et al., 2024). Such a configuration implies that the same voltage changes can induce more extended temporal delays in the generation of event signals in the low light condition. In Fig. 4(d), we display a real-world captured low-light event, where a temporal delay in the event data can be observed, leading to spatial trailing artifacts. Using such degraded event data for frame interpolation results in deformations in the interpolated frames due to the trailing effects (see Fig. 1).
Figure 4. Simplified circuit diagram for each pixel of the event camera. And the visualization of event trailing artifacts, captured by our EVS-RGB beam splitting imaging system.
3.2. Low Light Event-VFI
For common Event-VFI, the interpolation is modeled using Eqn. 1 and formulated as (Jiang et al., 2020; Sun et al., 2022; Lin et al., 2020):
@@ 3.2. Low Light Event-VFI
In recent methods (Tulyakov et al., 2021, 2022; Kim et al., 2023), to recover an intermediate RGB frame $I_{t+\Delta t}$ between $I_{t}$ and $I_{t+1}$, information from both the forward (from $I_{t}$ to $I_{t+\Delta t}$) and the backward (from $I_{t+1}$ to $I_{t+\Delta t}$) estimation are utilized. Thus process of low light Event-VFI in Eqn. 6 is solved as:
(7) $\hat{I}_{t+\Delta t}=\Phi(I_{t},I_{t+1},E_{t,t+\Delta t},E_{t+\Delta t,t+1},c,\alpha),$
where $\Phi$ is the interpolation algorithm.